OpenStudy (anonymous):

Write an equation and solve. Show the equation, your work in solving the equation, and the solution in your answer. Manuel went shopping and bought 3 pairs of pants. He also spent five dollars on a pair of socks. When he got home, he realized that he had spent a total of seventy-one dollars. What was the cost of each pair of pants?

OpenStudy (anonymous):

@texaschic101

OpenStudy (texaschic101):

3p + 5 = 71 3p = 71 - 5 3p = 66 p = 22 well...he spent $22 for all 3 pairs of pants, but I am not sure about how much he paid for each pair.....maybe, if the pants all cost the same price, then 22/3 = $7.33 a piece ??

OpenStudy (anonymous):

well 22/3 is 7.33 a piece so then I guess yeah

OpenStudy (texaschic101):

I hope thats right :)